OpenStudy (anonymous):

he formula for degrees Celsius is:5/9 C = (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 25 degrees Celsius is.

OpenStudy (andriod09):

tell me what do you know about the problem right now, as in what could you put down on a piece of paper?

OpenStudy (anonymous):

i have no idea this really confuses me

OpenStudy (andriod09):

okay. so then, tell me, from what you see, what do you think would be the first thing you should do?

OpenStudy (anonymous):

i dont know

OpenStudy (andriod09):

heres a hint, the first thing you should always do, is write down the formula. which in this case would be:

OpenStudy (anonymous):

5/9 C = (F − 32)

OpenStudy (andriod09):

okay, so now fill in the formula. \[\frac {5} {9} C=(F-32)\]

OpenStudy (andriod09):

do you have the formula filled in yet?

OpenStudy (anonymous):

idk how to

OpenStudy (andriod09):

what do ytou have that represents F? F=25, right?

OpenStudy (anonymous):

no

OpenStudy (andriod09):

then tell me, what is F and what is C?

OpenStudy (anonymous):

c is 25

OpenStudy (andriod09):

oh. okay then, that's simple, the answer is about 77. and just fill in the equation. its simple, ill start you out: \[\frac {5} {9}C = (F − 32)\]\[\frac {5} {9}25- (F-32)\] you can take it from there i hope